# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table 14. Identified monitoring challenge and solution for containerized applications in multi-cloud environment
ID Challenge Solution
Category 1: Performance Monitoring Consistency and Variability
(Almeida et al., 2014) Performance Monitoring Performance Variation Monitoring
(Jha et al., 2019) Benchmarking Performance Variation and Cost Efficiency Monitoring SDBO Optimization
(Gupta et al., 2021) Website Performance and Availability Monitoring Kubernetes Monitoring
(Pandey et al., 2022) User Experience Variability Monitoring CSP QoS Metrics Collection
(Dreibholz et al., 2019) Independent Performance Metrics Monitoring Performance Evaluations
(Bhamare et al., 2015) Comprehensive Network Performance Measurement Monitoring Server Selection Scheme
(Rahman and Lama, 2019) Microservice Performance Degradation Monitoring Tail Latency Prediction
(Struhár et al., 2024) Container Performance Metrics Tracking CLM and OSLM for Resource Adjustments
(Horchulhack et al., 2024) Container Performance and Resource Usage Monitoring Metric and Container Monitoring Tools
(Hossain et al., 2023) Microservice Performance Monitoring Challenge Prometheus, Grafana, Kubernetes for Monitoring
(Ramesh et al., 2023) Cross-Cloud Service Performance Monitoring FaaS and Storage Services for AI Monitoring
(Nguyen and Chien, 2023) Stream Processing Performance Monitoring RBAM for Real-Time Performance Insights
(Kataru et al., 2023) AWS Application Performance Monitoring AWS CloudWatch for Performance Monitoring
(Alonso et al., 2023a) Multi-Cloud Performance Visibility Issue Monitoring Tools for Multi-Cloud Visibility
(Harzenetter et al., 2023) Cross-Technology Performance Monitoring Issue TOSCA-based Workflows for Monitoring Applications
(Huang and Wang, 2023) Industrial Robot System Monitoring Challenge Real-Time System Performance Tracking Tools
Category 2: Resource and Infrastructure Management Monitoring
(Gattobigio et al., 2022) Network Communication Monitoring Service Mesh Implementation
(Verdugo et al., 2017) System Image and Configuration Monitoring Docker Server Deployment
(Cui et al., 2016) Vcluster and Framework Management Monitoring LXC Implementation
(Pahl, 2015) Multi-Cloud Resource Utilization Monitoring Prometheus and Grafana Monitoring
(Buyya and Barreto, 2015) Dynamic Resource Management Monitoring Aneka Platform
(da Silva et al., 2023) Cost and Resource Usage Monitoring COSTA for Cost and Resource Monitoring
(Dogani et al., 2023) Real-Time Dynamic Resource Monitoring Prometheus for Real-Time Monitoring
(Wong et al., 2023) Resource Usage and Performance Issue Prometheus for Real-Time Data Collection
(Marques et al., 2024) Real-Time Container Monitoring Prometheus with Forecasting for Resource Management
(Santos et al., 2023) Distributed Cloud Resource Monitoring Issue AWS CloudWatch and Azure Monitor for Insights
(Nagarajan et al., 2023) Resource Allocation and Energy Monitoring Issue Tools for Resource, Energy, and SLA Monitoring
Category 3: Security, Compliance, and Trust Management Monitoring
(Ramos et al., 2021) Provider Host Information Access Monitoring ML Monitoring
(Rios et al., 2017) Continuous Security Control Monitoring MUSA Platform
(Casola et al., 2017) Multi-Cloud Security Compliance Monitoring MUSA Security Compliance
(Acquaviva et al., 2017) Service Continuity and Data Privacy Monitoring NoMISHAP Implementation
(Wahab et al., 2016) Trust Relationship and Network Contact Monitoring Trust Bootstrapping
(Fugkeaw, 2023) Access and Authentication Monitoring Issue Blockchain for Secure Access Event Logging
(Wang et al., 2023) Malware Behavior and System Monitoring DockerWatch for Container Behavior Monitoring
(Neeraj et al., 2023) Log Data Integrity Monitoring Issue Blockchain-based Logs for SLA Monitoring
Category 4: Application Optimization and Adaptability Monitoring
(Ferry et al., 2014) Adaptive System Design and State Monitoring Cloud Modelling Framework
(Lorenzo et al., 2019) Partial Download Execution Monitoring FogDocker Implementation
(Wurster et al., 2018) Event-Driven Application Behavior Monitoring TOSCA Event Modeling
(Guerrero et al., 2018a) Microservice Optimization Monitoring MiCADO Optimization
(Bou Ghantous and Gill, 2021) Real-Time Application Monitoring DRA Framework
(Sofia et al., 2023) Container Application Adaptability Monitoring Advanced Monitoring with Orchestration Integration
(Petrovski and Gusev, 2024) Serverless and Containerized Monitoring Challenge AWS Tools for Performance Monitoring
Category 5: System Complexity and Standardization Monitoring
(Sabbioni et al., 2020) Proprietary Monitoring Solution Standardization Pervasive Monitoring
(Gao et al., 2017) Kernel Resource Isolation Mechanism Monitoring Leakage Defense
(Alonso et al., 2019) Multi-Cloud Application Monitoring Complexity DECIDE DevOps Expansion
(Barna et al., 2017) Monitoring Agent Coordination Autonomic Management
(Sami and Mourad, 2020) Heterogeneous Fog Resource Management Monitoring Kubernetes Utility Architecture
(Baresi et al., 2016) Application Topology and Dependency Monitoring MicroCloud TOSCA Library
(Trihinas et al., 2015) Elastic Cloud Service Monitoring Research JCatascopia Monitoring
(Fadda et al., 2019) Microservice Monitoring Diversity QoM Definition
(Harwalkar et al., 2019b) Compute and Network Resource Utilization Monitoring Time Series Resource Utilization
(Ahmad and Aujla, 2023) Real-Time Data Operations Monitoring Smart Contracts for GDPR Monitoring
(Centofanti et al., 2024) Power Consumption Monitoring Issue Prometheus with Power Monitoring Integration
(Tiwari and Sharma, 2023) Serverless Function Monitoring Complexity CLI Tool for Function Monitoring
Category 6: Multi-cloud Coordination and Integration Monitoring
(Sitaram et al., 2018) Integrated Multi-Cloud Management Monitoring Gnocchi Integration
(Mfula and Nurminen, 2018) Multi-Cloud Monitoring Challenges Self-Healing Techniques
(Kritikos et al., 2019) Cross-Level Monitoring Improvement Business Process Monitoring
(Lee and Nam, 2023) Real-Time Network Policy Monitoring Challenge KUNERVA for Real-Time Network Policy Monitoring
(Sofia et al., 2024) Decentralized System Observability Issue CODECO for Orchestration and Data Observability
Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment. Out of the 121 reviewed studies, 9 specifically deal with these concerns. One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns. Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications. Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment. This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management. The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring. The NoMISHAP Implementation is proposed to handle these monitoring issues. Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification. Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities. For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
@@
[…]
3.9. Tools and Frameworks (RQ6)
@@
In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in (Waseem et al., 2024)). Below, we provide an overview of each category, along with key tools and frameworks, highlighting their functionality, adoption, and distinctive features.
@@
Multi-cloud and Hybrid Cloud Platforms: are designed to support organizations in deploying and managing applications across different cloud providers smoothly. In this category, we identify and organize tools like Cloudify and CMP2, which offer end-to-end multi-cloud orchestration capabilities, along with platforms such as Cloudcheckr and MistIO, which provide monitoring and cost management across various cloud services. Similarly, AWS SDK and the CloudSME Simulation Platform (CSSP) are widely adopted for interacting with cloud APIs and simulating cloud environments, respectively. Moreover, hybrid platforms like OpenStack and VMware vSphere offer a blend of public and private cloud deployment options, which are essential for organizations pursuing a hybrid cloud strategy. This category highlights the tools that enable integration, scalability, and cost management for applications running across multiple clouds.
Networking, Security, and Access Management Tools: This category represents the tools that help to minimize security risks and ensure connectivity across containerized workloads in multi-cloud systems. As multi-cloud deployments become increasingly complex, advanced networking tools like Open vSwitch (OvS) enable developers to manage virtual networks across various cloud platforms. Security tools such as Seccomp-BPF and Linux Security Modules (LSM) provide critical protection to containers by enforcing strict access controls and sandboxing. Identity management solutions, such as Host Identity Protocol (HIP) and access control policies, ensure proper authentication and authorization for users and services in the cloud.
Monitoring, Management, and Automation Tools: Tools in this category provide the foundation for a scalable and automated approach to managing distributed applications by ensuring that workloads are continuously optimized and monitored on cloud platforms. Tools like Prometheus and Grafana offer powerful visualization capabilities for monitoring CPU usage, memory consumption, and network activity, aiding system administrators. With Terraform and Puppet, it is possible to automate infrastructure management, enabling developers to define infrastructure as code and automate provisioning across different cloud environments. Other tools, such as Ceilometer and Gnocchi, are responsible for gathering telemetry data in large OpenStack deployments.
@@
Table 15. Tools and frameworks for containerized applications in multi-cloud environment
@@
Container Orchestration and Management Kubernetes (Pai and Kumar, 2021; Wong et al., 2019; Gupta et al., 2021; Bolivar et al., 2018; del Castillo et al., 2013; Rodigari et al., 2021; Rahman and Lama, 2019; Pellegrini et al., 2017; Quint and Kratzke, 2018; Barletta et al., 2024; Struhár et al., 2024; Barati et al., 2023; Horchulhack et al., 2024; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Pandi et al., 2023; Sofia et al., 2024; Lee and Nam, 2023; Sofia et al., 2023; Nguyen and Chien, 2023; Tiwari and Sharma, 2023; Santos et al., 2023; Roy and Kalafatis, 2023; Kataru et al., 2023; Benmerar et al., 2023; Alonso et al., 2023a; Harzenetter et al., 2023; Huang and Wang, 2023; Wang et al., 2023; Nagarajan et al., 2023), Docker (Pham et al., 2015; Naik, 2016a; Ouyang et al., 2022; Slominski et al., 2015; Pai and Kumar, 2021; Gao et al., 2017; Alonso et al., 2019; Barna et al., 2017; Gupta et al., 2021; Sheridan et al., 2017; Mittal and Risco-Martín, 2017; Al-shammari and Alwan, 2018; Carpio et al., 2020; Bolivar et al., 2018; Rodigari et al., 2021; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ranjbar et al., 2017; Serhiienko and Spillner, 2018; Quint and Kratzke, 2018), Docker Swarm (Pai and Kumar, 2021; Wong et al., 2019; Altran et al., 2022; Rodigari et al., 2021; Quint and Kratzke, 2018), Mesos (Cui et al., 2016; Wong et al., 2019; Rodigari et al., 2021; Quint and Kratzke, 2018), LXC (Pahl, 2015; Gao et al., 2017; Gupta et al., 2021), OverlayFS (Al-shammari and Alwan, 2018), runC (Bélair et al., 2019), Container Technology (Wang et al., 2020), Docklet (Cui et al., 2016)
@@
Multi-cloud and Hybrid Cloud Platforms multi-cloud environment (Shi et al., 2020; Fadda et al., 2019; Di Modica et al., 2019; Tang, 2021), Cloudify (Zeck and Bouroudjian, 2017), CloudBroker Platform (Kiss, 2018), Cloudcheckr (Serhiienko and Spillner, 2018), MistIO (Serhiienko and Spillner, 2018), ManageIQ (Tang, 2021; Serhiienko and Spillner, 2018), CMP2 (Serhiienko and Spillner, 2018), AWS SDK (Graupner et al., 2015a), CloudSME Simulation Platform (CSSP) (Kiss, 2018), Vagrant (Pham et al., 2015), AWS OpsWorks (Pham et al., 2015), Universal Compute Xchange (UCX) (Zeck and Bouroudjian, 2017), Liferay portlets (Kiss, 2018), Cloud Service Providers (CSPs) (Graupner et al., 2015a), AWS Data Pipeline (Sitaram et al., 2018), GENI (Pandey et al., 2022), Google Cloud Platform (Pandey et al., 2022; Acquaviva et al., 2017; Tang, 2021; Quint and Kratzke, 2018), AWS EC2 (He et al., 2022; Buyya and Barreto, 2015; Pai and Kumar, 2021), Alibaba cloud (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), Tencent (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), OpenStack (Cui et al., 2016; Dreibholz et al., 2019), VMware vSphere (Mfula and Nurminen, 2018), Aliyun ECS (Hua et al., 2020), Baidu BCE (Hua et al., 2020), Tencent CW (Hua et al., 2020), JD VW (Hua et al., 2020), GoGrid (Buyya and Barreto, 2015)
@@
Networking, Security, and Access Management Open vSwitch (OvS) (Alaluna et al., 2017), Cloud networks (Ranjbar et al., 2017), Host Identity Protocol (HIP) (Ranjbar et al., 2017), Gateways (Harwalkar et al., 2019a), Pledge (Bélair et al., 2019), Seccomp-BPF (Bélair et al., 2019), Landlock LSM (Bélair et al., 2019), Linux Security Modules (LSM) (Bélair et al., 2019), MUSA Security Assurance Platform (Rios et al., 2017), Firewall Deployment (Sheridan et al., 2017), Access control policies (Gao et al., 2017), Attribute-Based Encryption and Signature (Shen et al., 2017), ABE/ABS cryptographic model (Shen et al., 2017)
@@
Monitoring, Management, and Automation Prometheus (Wong et al., 2019; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Grafana (Alonso et al., 2019; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Telegraf (Alonso et al., 2019), AWS CloudWatch (Alhamazani et al., 2014), cAdvisor (Wong et al., 2019), Ceilometer (del Castillo et al., 2013; Sitaram et al., 2018), Gnocchi (Sitaram et al., 2018), ELK stack (Mfula and Nurminen, 2018), Terraform (Alonso et al., 2019; Gupta et al., 2021; da Silva et al., 2023; Pandi et al., 2023; Harzenetter et al., 2023), CloudBees (Pham et al., 2015), Jenkins (Pandey et al., 2022; Solayman and Qasha, 2023), Puppet (Pham et al., 2015), Chef (Pham et al., 2015; Sheridan et al., 2017), GitHub (Sheridan et al., 2017), Shell scripts (Rios et al., 2017), Maven (Pham et al., 2015), NPM (Pham et al., 2015), Consul (Pham et al., 2015)
@@
Development, APIs, and Cloud Storage Public cloud storage (Harwalkar et al., 2019a), Private cloud storage (Harwalkar et al., 2019a), Hybrid cloud storage (Harwalkar et al., 2019a), Google Cloud Storage (Ramesh et al., 2023), Azure Blob Storage (Ramesh et al., 2023), REST API (Harwalkar et al., 2019a), YAML (Serhiienko and Spillner, 2018), OpenVAS Vulnerability Scanner (Sheridan et al., 2017), DAX format (Tang, 2021), Java (Sheridan et al., 2017; Rios et al., 2017; Carpio et al., 2020; Bucur et al., 2018; Quint and Kratzke, 2018; Challita et al., 2017), Go programming language (Al-shammari and Alwan, 2018; Lorenzo et al., 2019), Python (Carpio et al., 2020; Bucur et al., 2018; del Castillo et al., 2013; Sitaram et al., 2018; Serhiienko and Spillner, 2018), Ruby (Bucur et al., 2018), PHP (Bucur et al., 2018), C# (Bucur et al., 2018), Cloud Modelling Framework (CLOUMDF) (Ferry et al., 2014), Cloud Modelling Language (CLOUDML) (Ferry et al., 2014; Challita et al., 2017), CAMEL modeling language (Casola et al., 2017; Quint and Kratzke, 2018; Kritikos et al., 2019), Eclipse IDE (Alonso et al., 2019), MCSLA editor (Alonso et al., 2019)
[…]
4. Discussion
4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)

### v3

@@ 3.8 Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment. Out of the 121 reviewed studies, 9 specifically deal with these concerns. One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns.
Table 14: Identified monitoring challenges and solutions for containerized applications in multi-cloud environment
      ID        Challenge        Solution
      Category 1: Performance Monitoring Consistency and Variability
      [95]        Performance Monitoring        Performance Variation Monitoring
      [96]        Benchmarking Performance Variation and Cost Efficiency Monitoring        SDBO Optimization
      [103]        Website Performance and Availability Monitoring        Kubernetes Monitoring
      [67]        User Experience Variability Monitoring        CSP QoS Metrics Collection
      [122]        Independent Performance Metrics Monitoring        Performance Evaluations
      [47]        Comprehensive Network Performance Measurement Monitoring        Server Selection Scheme
      [53]        Microservice Performance Degradation Monitoring        Tail Latency Prediction
      [137]        Container Performance Metrics Tracking        CLM and OSLM for Resource Adjustments
      [87]        Container Performance and Resource Usage Monitoring        Metric and Container Monitoring Tools
      [37]        Microservice Performance Monitoring Challenge        Prometheus, Grafana, Kubernetes for Monitoring
      [92]        Cross-Cloud Service Performance Monitoring        FaaS and Storage Services for AI Monitoring
      [119]        Stream Processing Performance Monitoring        RBAM for Real-Time Performance Insights
      [134]        AWS Application Performance Monitoring        AWS CloudWatch for Performance Monitoring
      [33]        Multi-Cloud Performance Visibility Issue        Monitoring Tools for Multi-Cloud Visibility
      [141]        Cross-Technology Performance Monitoring Issue        TOSCA-based Workflows for Monitoring Applications
      [49]        Industrial Robot System Monitoring Challenge        Real-Time System Performance Tracking Tools
      Category 2: Resource and Infrastructure Management Monitoring
      [34]        Network Communication Monitoring        Service Mesh Implementation
      [120]        System Image and Configuration Monitoring        Docker Server Deployment
      [126]        Vcluster and Framework Management Monitoring        LXC Implementation
      [60]        Multi-Cloud Resource Utilization Monitoring        Prometheus and Grafana Monitoring
      [98]        Dynamic Resource Management Monitoring        Aneka Platform
      [115]        Cost and Resource Usage Monitoring        COSTA for Cost and Resource Monitoring
      [116]        Real-Time Dynamic Resource Monitoring        Prometheus for Real-Time Monitoring
      [136]        Resource Usage and Performance Issue        Prometheus for Real-Time Data Collection
      [88]        Real-Time Container Monitoring        Prometheus with Forecasting for Resource Management
      [125]        Distributed Cloud Resource Monitoring Issue        AWS CloudWatch and Azure Monitor for Insights
      [43]        Resource Allocation and Energy Monitoring Issue        Tools for Resource, Energy, and SLA Monitoring
      Category 3: Security, Compliance, and Trust Management Monitoring
      [132]        Provider Host Information Access Monitoring        ML Monitoring
      [56]        Continuous Security Control Monitoring        MUSA Platform
      [127]        Multi-Cloud Security Compliance Monitoring        MUSA Security Compliance
      [146]        Service Continuity and Data Privacy Monitoring        NoMISHAP Implementation
      [72]        Trust Relationship and Network Contact Monitoring        Trust Bootstrapping
      [144]        Access and Authentication Monitoring Issue        Blockchain for Secure Access Event Logging
      [142]        Malware Behavior and System Monitoring        DockerWatch for Container Behavior Monitoring
      [147]        Log Data Integrity Monitoring Issue        Blockchain-based Logs for SLA Monitoring
      Category 4: Application Optimization and Adaptability Monitoring
      [89]        Adaptive System Design and State Monitoring        Cloud Modelling Framework
      [113]        Partial Download Execution Monitoring        FogDocker Implementation
      [140]        Event-Driven Application Behavior Monitoring        TOSCA Event Modeling
      [99]        Microservice Optimization Monitoring        MiCADO Optimization
      [90]        Real-Time Application Monitoring        DRA Framework
      [118]        Container Application Adaptability Monitoring        Advanced Monitoring with Orchestration Integration
      [100]        Serverless and Containerized Monitoring Challenge        AWS Tools for Performance Monitoring
      Category 5: System Complexity and Standardization Monitoring
      [65]        Proprietary Monitoring Solution Standardization        Pervasive Monitoring
      [70]        Kernel Resource Isolation Mechanism Monitoring        Leakage Defense
      [105]        Multi-Cloud Application Monitoring Complexity        DECIDE DevOps Expansion
      [52]        Monitoring Agent Coordination        Autonomic Management
      [29]        Heterogeneous Fog Resource Management Monitoring        Kubernetes Utility Architecture
      [31]        Application Topology and Dependency Monitoring        MicroCloud TOSCA Library
      [84]        Elastic Cloud Service Monitoring Research        JCatascopia Monitoring
      [85]        Microservice Monitoring Diversity        QoM Definition
      [78]        Compute and Network Resource Utilization Monitoring        Time Series Resource Utilization
      [143]        Real-Time Data Operations Monitoring        Smart Contracts for GDPR Monitoring
      [139]        Power Consumption Monitoring Issue        Prometheus with Power Monitoring Integration
      [131]        Serverless Function Monitoring Complexity        CLI Tool for Function Monitoring
      Category 6: Multi-cloud Coordination and Integration Monitoring
      [81]        Integrated Multi-Cloud Management Monitoring        Gnocchi Integration
      [42]        Multi-Cloud Monitoring Challenges        Self-Healing Techniques
      [76]        Cross-Level Monitoring Improvement        Business Process Monitoring
      [145]        Real-Time Network Policy Monitoring Challenge        KUNERVA for Real-Time Network Policy Monitoring
      [138]        Decentralized System Observability Issue        CODECO for Orchestration and Data Observability
Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications. Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment. This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management. The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring. The NoMISHAP Implementation is proposed to handle these monitoring issues. Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification. Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities. For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
Takeaway 11: Reviewed studies reveal that monitoring containerized applications in multi-cloud environments introduces unique challenges, such as inconsistent performance metrics across cloud providers, real-time trust and security event tracking, and fragmented observability due to container orchestration spanning heterogeneous infrastructures. These challenges are addressed through specialized solutions like cross-cloud Prometheus-Grafana stacks, blockchain-based SLA logging, and service mesh integrations, reflecting the need for tightly coupled monitoring mechanisms tailored specifically to the hybrid nature of containerized multi-cloud deployments.
3.9 Tools and Frameworks (RQ6)
@@
In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in [11]). Below, we provide an overview of each category, along with key tools and frameworks, highlighting their functionality, adoption, and distinctive features.
@@
Multi-cloud and Hybrid Cloud Platforms are designed to support organizations in deploying and managing applications across different cloud providers smoothly. In this category, we identify and organize tools like Cloudify and CMP2, which offer end-to-end multi-cloud orchestration capabilities, along with platforms such as Cloudcheckr and MistIO, which provide monitoring and cost management across various cloud services. Similarly, AWS SDK and the CloudSME Simulation Platform (CSSP) are widely adopted for interacting with cloud APIs and simulating cloud environments, respectively. Moreover, hybrid platforms like OpenStack and VMware vSphere offer a blend of public and private cloud deployment options, which are essential for organizations pursuing a hybrid cloud strategy. This category highlights the tools that enable integration, scalability, and cost management for applications running across multiple clouds.
Networking, Security, and Access Management Tools: This category represents the tools that help to minimize security risks and ensure connectivity across containerized workloads in multi-cloud systems.
Advanced networking tools like Open vSwitch (OvS) allow developers to manage virtual networks across multiple cloud platforms as multi-cloud deployments grow more complicated. By implementing stringent access controls and sandboxing, security tools like Seccomp-BPF and Linux Security Modules (LSM) offer vital container protection. Proper authentication and authorization for cloud users and services are ensured by identity management solutions like Host Identity Protocol (HIP) and access control policies.
Monitoring, Management, and Automation Tools: Tools in this category help manage large-scale cloud applications in an easy and automated way. They make sure everything runs smoothly by constantly checking and improving how resources are used. For example, Prometheus and Grafana help system administrators by showing important information like CPU use, memory load, and network traffic in clear charts and dashboards. Tools like Terraform and Puppet make it possible to write and reuse code to set up servers and other infrastructure automatically across different cloud services. This saves time and reduces errors. In big systems using OpenStack, tools such as Ceilometer and Gnocchi collect data about how the system is working, which helps with tracking performance and solving issues.
@@
Table 15: Tools and frameworks for containerized applications in multi-cloud environment
@@
       Container Orchestration and Management        Kubernetes [59, 111, 103, 40, 107, 35, 53, 73, 50, 82, 137, 135, 87, 116, 136, 88, 110, 36, 143, 139, 37, 130, 91, 117, 138, 145, 118, 119, 131, 125, 46, 134, 104, 33, 141, 49, 142, 43], Docker [86, 148, 121, 44, 59, 70, 105, 52, 103, 55, 106, 112, 30, 40, 35, 53, 42, 68, 45, 50], Docker Swarm [59, 111, 114, 35, 50], Mesos [126, 111, 35, 50], LXC [60, 70, 103], OverlayFS [112], runC [71], Container Technology [66], Docklet [126]
@@
       Multi-cloud and Hybrid Cloud Platforms        multi-cloud environment [97, 85, 129, 41], Cloudify [62], CloudBroker Platform [63], Cloudcheckr [45], MistIO [45], ManageIQ [41, 45], CMP2 [45], AWS SDK [57], CloudSME Simulation Platform (CSSP) [63], Vagrant [86], AWS OpsWorks [86], Universal Compute Xchange (UCX) [62], Liferay portlets [63], Cloud Service Providers (CSPs) [57], AWS Data Pipeline [81], GENI [67], Google Cloud Platform [67, 146, 41, 50], AWS EC2 [54, 98, 59], Alibaba cloud [62, 50], Tencent [62, 50], OpenStack [126, 122], VMware vSphere [42], Aliyun ECS [93], Baidu BCE [93], Tencent CW [93], JD VW [93], GoGrid [98]
@@
       Networking, Security, and Access Management        Open vSwitch (OvS) [124], Cloud networks [68], Host Identity Protocol (HIP) [68], Gateways [108], Pledge [71], Seccomp-BPF [71], Landlock LSM [71], Linux Security Modules (LSM) [71], MUSA Security Assurance Platform [56], Firewall Deployment [55], Access control policies [70], Attribute-Based Encryption and Signature [109], ABE/ABS cryptographic model [109]
@@
       Monitoring, Management, and Automation        Prometheus [111, 116, 136, 88, 110, 36, 143, 139, 37, 130, 91, 46], Grafana [105, 37, 130, 91, 46], Telegraf [105], AWS CloudWatch [83], cAdvisor [111], Ceilometer [107, 81], Gnocchi [81], ELK stack [42], Terraform [105, 103, 115, 117, 141], CloudBees [86], Jenkins [67, 38], Puppet [86], Chef [86, 55], GitHub [55], Shell scripts [56], Maven [86], NPM [86], Consul [86]
@@
       Development, APIs, and Cloud Storage        Public cloud storage [108], Private cloud storage [108], Hybrid cloud storage [108], Google Cloud Storage [92], Azure Blob Storage [92], REST API [108], YAML [45], OpenVAS Vulnerability Scanner [55], DAX format [41], Java [55, 56, 30, 133, 50, 94], Go programming language [112, 113], Python [30, 133, 107, 81, 45], Ruby [133], PHP [133], C# [133], Cloud Modelling Framework (CLOUMDF) [89], Cloud Modelling Language (CLOUDML) [89, 94], CAMEL modeling language [127, 50, 76], Eclipse IDE [105], MCSLA editor [105]
Takeaway 12: The reviewed studies reveal that tools and frameworks commonly used in general container or cloud settings (e.g., Kubernetes, Docker, Terraform, Prometheus) require specific adaptations or configurations to operate effectively in containerized multi-cloud environments. This includes challenges such as consistent orchestration across heterogeneous cloud APIs, maintaining observability across isolated infrastructures, and enabling secure, policy-compliant automation at scale. Moreover, hybrid-enabling tools like OpenStack, Cloudify, and MistIO emerged as critical for unifying management across cloud boundaries — highlighting how the combination of containerization and multi-cloud imposes unique operational and integration demands not encountered when using either paradigm alone.
4 Discussion
The following discussion synthesizes key findings for each research question and their implications for research and practice. These implications also highlight open problems and unresolved challenges that require further research attention, particularly those emerging from the intersection of containerization and multi-cloud environments.
4.1 Containers in Multi-Cloud: Roles and Strategies (RQ1)